Determine all the relative minimum and maximum values, and saddle points of the function g defined by g(x,y) = x3 - 3x +3xy2

1
Expert's answer
2022-05-12T18:25:15-0400

"f_x=3x^2-3+3y^2"

"f_{xx}=6x"

"f_y=6xy"

"f_{yy}=6x"

"f_{xy}=6y"

Critical points

"f_x=3x^2-3+3y^2=0"

"f_y=6xy=0"

We can set them equal to each other

"3x^2-3+3y^2=6xy"

"x^2-2xy+y^2=1"

"(x-y)^2=1"

x-y=1

y=x-1....[i]

Solving fy=0 gives


Either x=0 , or y=0

When x=0,y=−1⟹(x,y)=(0,−1)


When y=0,x=1⟹(x,y)=(1,0)

We know


D=fxxfyy−[fxy]2


We require D<0 for the point (x,y) to be a saddle point


When (x,y)=(0,−1)

"D=6(0)6(0)-(6(-1))^2=-36"

straightforward result for a confirmed saddle point.


When (x,y)=(1,0)

"D=6(1)6(1)-(6(0))^2=36>0"

This is not saddle point. As D>0 and fxx>0, so it is relative minimum.

Conclusion: The saddle point is (x,y)=(0,−1), relative minimum is (1,0)